OpenStudy (anonymous):

Robin chose a slice from a pizza with a 20" diameter. The central angle of her slice of pizza is measured 60. What was the area of Robin's pizza?

OpenStudy (danjs):

60 degrees is 1/6 th of the whole circle. Find the area of a circle with a 20" diameter, and then take a sixth of that for the area of a 60 degree slice

OpenStudy (anonymous):

@DanJS so is the answer 209.44?

OpenStudy (danjs):

\[Pizza = \pi *10^2 = 100\pi\] \[60~degree~slice = \frac{ whole ~pizza }{ 6 } = \frac{ 100\pi }{ 6 } ~In^2\]

OpenStudy (danjs):

remember the area of a circle is pi times the radius squared, not the diameter squared

OpenStudy (anonymous):

Thanks!! :) @DanJS

OpenStudy (danjs):

yw